Please help me solve this: use function notation to write the equation of the line passing through (-2,-3) and (2,5)

Answer by stanbon(75887)

write the equation of the line passing through (-2,-3) and (2,5)
----------------------
slope = (5--3)/(2--2) = 8/4 = 2
--------------------------------------
Line form is y = mx + b
Substitute and solve for "b":
5 = 2(2) + b
b = 1
-------
Equation:
y = 2x + 1
=========================
Cheers,
Stan H.